# Discovering dense hydrogen solid at 1200K with deep variational free energy approach

### v1

@@ I Introduction
In spite of extensive computational studies of the liquid-liquid transition in the hydrogen system, its underlying physics remains unclear. This is partly due to that traditional methods, such as density functional theory (DFT) or ab-initio molecular dynamics, are limited by their accuracy [15, 16, 17]. More accurate yet computationally expensive methods, such as quantum Monte Carlo, also face challenges in unbiasedly sampling the proton degrees of freedom [18, 19, 20, 21, 22]. Moreover, most of these approaches suffer from high computational costs, which limit simulations to relatively small systems or short timescales. Recent efforts based on machine-learned force fields [23] have enabled simulations at much larger sizes and longer timescales. However, the possible trade-off in accuracy during the supervised training [24] may limit their capacity to fully capture the complex behavior of the underlying physical processes. For example, the latest molecular dynamics studies based on the machine learning force field [25, 26] show qualitatively different results compared to earlier works [23]. In this context, it is noteworthy to mention recent findings on the solid-state phase around 1200 K [27, 28] obtained using machine learning force fields trained with quantum Monte Carlo (QMC) data, which sharply contrast with previous DFT calculations and experimental observations.
@@ II.1 Variational free energy framework
$p(\bm{X})=\frac{e^{-E_{\mathrm{base}}(\bm{Z})/k_{B}T}}{\mathcal{Z}_{\mathrm{base}}}\left|\det{\left(\frac{\partial\bm{Z}}{\partial\bm{X}}\right)}\right|\,,$ (3)
where $E_{\mathrm{base}}$ is the potential energy surface (PES) provided by the base model, $\bm{Z}$ are a set of collective variables that are transformed from the proton coordinates $\bm{X}$ via a neural network with a non-degenerate Jacobian. By parametrizing the proton Boltzmann distribution in this way, the normalizing flow model learns to correct the potential bias in the base PES. The partition function of the base model, $\mathcal{Z}_{\mathrm{base}}$, acts as a constant factor across different parameters and does not affect the optimization process. The Jacobian determinant of the transformation in Eq. (3) ensures that the normalization factor of the proton distribution remains independent of the flow model parameters. With a perfectly optimized flow model, the probability distribution will converge to the target distribution $p(\bm{X})\propto e^{-E(\bm{X})/k_{B}T}$.
@@
To better leverage prior knowledge about the hydrogen system within the optimization framework, we design a two-stage optimization scheme as shown in Figure 1. First, proton configurations are generated through molecular dynamics sampling using DeePMD, a machine learning-based potential energy model [37]. Next, we pretrain the electronic wavefunction on the ensemble of proton configurations [38, 39]. Finally, the wavefunction and the flow model—where the DeePMD model serves as the base distribution [Eq. (3)]—are jointly optimized. This approach ensures that the variational free energy calculation serves as a fine-tuning stage that corrects possible biases in the DeePMD model for the proton distribution. In the following subsection, we will discuss general considerations for designing neural network architectures for variational free energy optimization of liquid hydrogen.
@@ II.2 Neural networks
For for the electron wavefunction ansatz, we design a modified Ferminet [40] / Deepsolid [41] network architecture to satisfy the required symmetry constraints in the liquid phase. The translation invariance of the electron wavefunction under translations of both the electrons and protons requires $\Psi(\bm{R};\bm{X})=\Psi(\bm{R}+\bm{s};\bm{X}+\bm{s})$, where $\bm{s}$ is any constant vector in $\mathbb{R}^{3}$. Additionally, we impose the special twisted boundary conditions (TBC) [42] on the electronic degrees of freedom to mitigate finite-size effects. Under these conditions, wrapping an electron around the simulation cell introduces an additive phase change to the wavefunction, while the wavefunction remains periodic with respect to the proton degrees of freedom
@@
with $\bm{n}=n_{1}\bm{a}_{1}+n_{2}\bm{a}_{2}+n_{3}\bm{a}_{3}$, $n_{\mu}\in\mathbb{Z}$ and $\mu=1,2,3$, and $\bm{w}$ the specified twist of the wavefunction. Furthermore, the wavefunction should be symmetric (anti-symmetric) with respect to the permutation of protons (electrons of the same spin):
@@ III.1 Benchmarks
In our framework, the wavefunction is pretrained using ensemble VMC training. In the pretraining procedure, the proton configurations are sampled using DeePMD simulations on a grid within the temperature-$r_{s}$ thermodynamic state space. The temperature ranges from 1000 to 2000 K in intervals of 200 K, while $r_{s}$ ranges from 1.35 to 1.50 in intervals of 0.02. A total of 6400 configurations are generated, and a random subset with 512 configurations is used for each training step. The batch size for electron configurations is set to 8.
@@
$r_{s}=1.40$ 1230.81 1231.16 1230.83
$r_{s}=1.44$ 1231.61 1231.63 1231.63
To test the transferability of our wavefunction, we benchmark its performance by comparing the ground-state energies predicted by the multi-configuration trained model against those of the single-configuration trained model. Table 1 reports the mean absolute error (MAE) calculated using three test configurations for $r_{s}=1.40$ and $r_{s}=1.44$ at three different temperatures. These tested thermodynamic states are not included in the training ensemble. In all cases, the MAE remains below 2 mHa per atom, demonstrating that our pre-trained wavefunction is able to reliably provides accurate energy estimates across an ensemble of configurations, thus satisfying the requirements for variational free energy training.
Figure 2: The (a) pressure and (b) entropy per particle versus density parameteter computed at $T=1200$K for $N=32$ hydrogen atoms. The variational free energy calculations reveal significant features due to the formation of a crystalline structure at low densities.
@@
Figure 4: (a) The radius distributuion function of protons computed at $T=1200$K and $r_{s}=1.43$ along with the one after quench. (b) A snapshot of the protons shown as white spheres. The red spheres denote the center of mass of hydrogen molecules. (c) The simulated XRD spectra for an ensemble of configurations obtained at 1200K.
@@ III.2 High-temperature solids
In this subsection, we present the results of variational free energy optimization for the dense hydrogen system at 1200 K, focusing on the emergence of crystalline order.
@@
Figure 2 shows the pressure and entropy as a function of the density parameter $r_{s}$ at $T=1200$K. The change in slope at $r_{s}=1.42$ in the pressure curve indicates a discontinuous phase transition. Since the value of the partition function of the base distribution enters the measurement of entropy (see Eq. (3)), we performed annealed importance sampling [43] on the DeePMD model to compute its free energy $-\ln\mathcal{Z}_{DP}$ in order to obtain the entropy of the system. The entropy curves shown in panel (b) provides further insight into the two states, with the sharp decrease in entropy at $r_{s}=1.42$ suggesting a transition to a more ordered state.
@@
To further analyze the phase transition, we plot the proton-proton radial distribution function (RDF) of the trained model alongside the DeePMD base distribution in Figure 3. Compared to the uniform base distribution employed in Ref. 29, the DeePMD model provides a more informative reference for the proton distribution. However, around the critical point at $r_{s}=1.42$, the variational free energy optimization results in noticeable deviations in the RDF. The transition into a molecular state is characterized by the emergence of the molecular peak at around $1.4~{}a_{B}$ as the $r_{s}$ value increases. Additionally, the presence of extra features around $2~{}a_{B}$ and $3~{}a_{B}$ suggests that the system may develop additional structures other than a molecular liquid.
@@
We perform a detailed analysis of the proton configurations for $r_{s}$ values larger than 1.41 to gain deeper insight into the molecular phase. Panel (a) of Figure 4 presents a snapshot of the sampled atomic configurations at $r_{s}=1.43$. Hydrogen molecules are identified by pairing the closest hydrogen atoms, with their centers of mass marked by red spheres. Notably, these red spheres do not follow a random liquid arrangement, but instead form an ordered structure reminiscent of a solid. The averaged simulated X-ray diffraction (XRD) spectra for an ensemble of hydrogen configurations at three different $r_{s}$ values are shown in Figure 4 (c). The XRD spectra of the structures do not show significant similarities to the $Fmmm$ structure reported in Ref. 27.
To further investigate the crystal symmetry of the discovered solid structure, we quench the atomic configurations by sampling from $p(\bm{X})^{\beta}$ with $\beta=20$. Figure  4 (b) shows the RDF of the original and quenched structures. As the system cools down, the features around $2~{}a_{B}$ and $3~{}a_{B}$ become more prominent. Symmetry analysis of the centers of mass of the hydrogen atoms in the quenched samples reveals that most of the configuraionts belong to the $R\bar{3}c$ space group. On the other hand, symmetry analysis of the hydrogen atom coordinates reveals predominantly $P1$ symmetry, indicating a lack of orientational order in the H2 molecules.
@@ IV Conclusion and Discussion
In this work, we observe evidence of a direct transition from atomic liquid to molecular crystal within a temperature regime believed to be relevant to the liquid-liquid transition in the dense hydrogen system. This finding arises from a variational free energy calculation with many-body precision. Moreover, our results demonstrate that using many-body precision energy, the flow model can transform the structureless liquid state of the base proton distribution into a solid-state solution. However, since the simulation is performed with a finite box containing 32 hydrogen atoms, finite-size effects may bias both the existence and the structure of the discovered solid states. Nevertheless, this finding aligns with recent discoveries of a possible solid-state phase in similar temperature and pressure regimes obtained through different computational methods and much larger simulation cells [27, 28], although with different detailed characterizations of the solid-state phase. Collectively, these works suggest that the liquid to solid transition may ocurr at a higher temperature than previously believed. Replacing the expensive wavefunction model with a cheaper yet accurate machine learning force field could allow larger scale calculations, potentially bypassing some of the limitations of this study. Furthermore, nuclear quantum effects can be incorporated into the calculation by generalizing the proton distribution to a variational density matrix, using the neural canonical transformation approach [44, 45].
@@ Acknowledgements.
We thank Guglielmo Mazzola, Yubo Yang, and Qi Zhang for useful discussions. This project is supported by the National Natural Science Foundation of China under Grants No. 12122103, No. 92270107, No. T2225018, No. 12188101, No. T2121001 and the Strategic Priority Research Program of Chinese Academy of Sciences under Grants No. XDB0500000 and No. XDB30000000.

### v2

# Deep variational free energy prediction of dense hydrogen solid at 1200K

@@ I Introduction
In spite of extensive computational studies of the liquid-liquid transition in the hydrogen system, its underlying physics remains unclear. This is partly due to that traditional methods, such as density functional theory (DFT) or ab-initio molecular dynamics, are limited by their accuracy [15, 16, 17]. More accurate methods like quantum Monte Carlo (QMC), while providing higher precision, also face challenges in achieving unbiased sampling of proton degrees of freedom [18, 19, 20, 21, 22]. Moreover, most of these approaches suffer from high computational costs, which limit simulations to relatively small systems or short timescales. Recent efforts based on machine-learned force fields [23] have enabled simulations at much larger sizes and longer timescales. However, the possible trade-off in accuracy during the supervised training [24] may limit their capacity to fully capture the complex behavior of the underlying physical processes. For example, the latest molecular dynamics studies based on the machine learning force field trained with QMC data [25, 26] find first-order atomic liquid to molecular liquid transition around 1200K, which is qualitatively different from the supercriticality behavior reported in Ref. 23. In this context, it is noteworthy to mention recent findings on the solid-state phase around 1200 K [27, 28] obtained using machine learning force fields trained with QMC data, which sharply contrast with all of previous calculations and experimental observations. The discrepancies between these various experimental and computational studies might be caused by multiple reasons. On the experimental side, due to the difficulty of performing measurements under high pressures, experiments don’t directly measure melting but instead detect anomalies in the Raman signal. Additionally, the temperature at which melting occurs is indirectly inferred. On the computational side, recent findings in Refs. 27, 28 indicate that the accuracy of the potential energy surface can qualitatively affect the predicted phases, highlighting the importance of incorporating precise many-body energies beyond the DFT level. These considerations motivate a careful and independent re-examination of dense hydrogen in the debated parameter regime.
@@ II.1 Variational free energy framework
$p(\bm{X})=p_{\text{base}}(\bm{Z})\left|\det{\left(\frac{\partial\bm{Z}}{\partial\bm{X}}\right)}\right|=\frac{e^{-E_{\mathrm{base}}(\bm{Z})/k_{B}T}}{\mathcal{Z}_{\mathrm{base}}}\left|\det{\left(\frac{\partial\bm{Z}}{\partial\bm{X}}\right)}\right|\,,$ (3)
where $E_{\mathrm{base}}$ is the potential energy surface (PES) provided by the base model, $\bm{Z}$ are a set of collective variables that are transformed from the proton coordinates $\bm{X}$ via a neural network with a non-degenerate Jacobian. By parametrizing the proton Boltzmann distribution in this way, the normalizing flow model learns to correct the potential bias in the base PES model. The Jacobian determinant of the transformation in Eq. (3) ensures that the normalization factor of the proton distribution remains independent of the flow model parameters. With a perfectly optimized flow model, the probability distribution will converge to the target distribution $p(\bm{X})\propto e^{-E(\bm{X})/k_{B}T}$.
@@
During the variational free energy training process, the objective function in Eq. (2) is optimized by adjusting the parameters of both the flow model [Eq. (3)] and the electron wavefunction network [Eq. (4)] via stochastic optimization. At each optimization step, we use Monte Carlo algorithm to draw samples of proton and electron coordinates from both models via ancestral sampling. To better leverage prior knowledge about the hydrogen system within the optimization framework, we design a two-stage optimization scheme as shown in Figure 1. The key idea is to train a machine-learning-based potential energy model—DeePMD [37]—and use it to facilitate the optimization of both the flow model and the electron wavefunction (see Appendix C for details of the DeePMD model). More specifically, in the pretraining stage, we optimize the electronic wavefunction on an ensemble of proton configurations [38, 39] generated through molecular dynamics sampling using DeePMD. This make sure the electron wavefunction is at a good starting point before the joint optimization. Next, the wavefunction and the flow model—where the DeePMD model serves as the base distribution in Eq. (3)—are jointly optimized during the variational training process. This approach ensures that the variational free energy calculation serves as a fine-tuning stage, which corrects possible biases in the DeePMD model for the proton distribution while preserving a fully variational framework and yielding a proton distribution consistent with QMC-level accuracy.
In the following subsection, we will discuss general considerations for designing neural network architectures for variational free energy optimization of liquid hydrogen.
@@ II.2 Neural networks
For the electron wavefunction ansatz, we design a modified Ferminet [40] / Deepsolid [41] network architecture to satisfy the required symmetry constraints in the liquid phase. The translation invariance of the electron wavefunction under translations of both the electrons and protons requires $\Psi(\bm{R};\bm{X})=\Psi(\bm{R}+\bm{s};\bm{X}+\bm{s})$, where $\bm{s}$ is any constant vector in $\mathbb{R}^{3}$. Additionally, we impose the special twisted boundary conditions (TBC) [42] on the electronic degrees of freedom to mitigate finite-size effects. Under these conditions, wrapping an electron around the simulation cell introduces an additive phase change to the wavefunction, while the wavefunction remains periodic with respect to the proton degrees of freedom
@@
with $\bm{n}=n_{1}\bm{a}_{1}+n_{2}\bm{a}_{2}+n_{3}\bm{a}_{3}$, $n_{\mu}\in\mathbb{Z}$ and $\mu=1,2,3$, and $\bm{w}$ the specified twist of the wavefunction. Furthermore, the wavefunction should be symmetric (anti-symmetric) with respect to the permutation of protons (electrons of the same spin) $\sigma$
@@ III.1 Benchmarks
$r_{s}=1.40$ 1230.8(1) 1231.2(5) 1230.8(2)
$r_{s}=1.44$ 1231.6(3) 1231.6(3) 1231.6(3)
First, to demonstrate the accuracy of our electron wavefunction neural network in single-configuration variational calculations, we benchmark it against the reference data for 64 hydrogen atoms from Ref. [43]. As shown in Appendix B, our wavefunction consistently achieves lower ground-state energies than the reference data across all thermodynamic states in single-configuration training tasks, including both molecular and atomic configurations.
Next, we benchmark the accuracy of the multi-configuration-trained wavefunction at the pretraining stage to test its transferability and ensure it is a good starting point for the variational free energy training. In the pretraining step, the wavefunction is pretrained using ensemble VMC training. Proton configurations are sampled using DeePMD simulations over a grid in the temperature-$r_{s}$ thermodynamic state space. Temperatures range from 1000K to 2000K in increments of 200K, while $r_{s}$ ranges from 1.35 to 1.50 in steps of 0.02. A total of 6400 configurations are generated, from which a random subset of 512 configurations is used at each training step. The batch size for electron configurations is set to 8 for each proton configuration. To test the transferability of our wavefunction, we benchmark its performance by comparing the ground-state energies predicted by the multi-configuration trained model against those of the single-configuration trained model. Table 1 reports the mean absolute error (MAE) calculated using three test configurations for $r_{s}=1.40$ and $r_{s}=1.44$ at three different temperatures. These tested thermodynamic states are not included in the training ensemble. In all cases, the MAE remains below 2 mHa per atom, demonstrating that our pre-trained wavefunction is able to reliably provides accurate energy estimates across an ensemble of configurations, thus satisfying the requirements for variational free energy training.
Figure 2: The (a) pressure and (b) entropy per particle versus density parameter computed at $T=1200$K for $N=32$ hydrogen atoms. The variational free energy calculations reveal significant features due to the formation of a crystalline structure at low densities.
@@
Figure 4: (a) The radial distribution function of protons computed at $T=1200$K and $r_{s}=1.43$ along with the one after quench. (b) A snapshot of the protons shown as white spheres. The red spheres denote the center of mass of hydrogen molecules. (c) The simulated XRD spectra for an ensemble of configurations obtained at 1200K.
@@ III.2 High-temperature solids
In this subsection, we present the results of variational free energy optimization for the dense hydrogen system at 1200K, with a focus on the emergence of crystalline order. Both the flow model and the electron wavefunction are trained within the variational free energy framework. At each optimization step, 512 proton configurations are sampled from the trained Boltzmann distribution, and 16 electron configurations are sampled for each proton configuration.
@@
Figure 2 shows the pressure and entropy as a function of the density parameter $r_{s}$ at $T=1200$K. The change in slope at $r_{s}=1.42$ in the pressure curve indicates a discontinuous phase transition. Since the partition function of the base distribution enters the entropy calculation (see Eq. (3)), we performed annealed importance sampling [44] on the DeePMD model to compute its free energy $-\ln\mathcal{Z}_{DP}$ in order to obtain the entropy of the system. The entropy curves shown in panel (b) provides further insight into the two states, with the sharp decrease in entropy at $r_{s}=1.42$ suggesting a transition to a more ordered state.
@@
To further analyze the phase transition, we plot the proton-proton radial distribution function (RDF) of the trained model alongside the DeePMD base distribution in Figure 3. Compared to the uniform base distribution employed in Ref. 29, the DeePMD model provides a more informative reference for the proton distribution. However, around the critical point at $r_{s}=1.42$, the variational free energy optimization leads to noticeable deviations in the radial distribution function (RDF). The transition into a molecular state is characterized by the emergence of the molecular peak at around $1.4~{}a_{B}$ as the $r_{s}$ value increases. Additionally, the presence of extra features around $2~{}a_{B}$ and $3~{}a_{B}$ suggests that the system may develop additional structures other than a molecular liquid. The convergence of the flow-based proton distributions is examined in Appendix D. The results shown in Fig. 6 indicate that as the flow model is trained, the proton distribution converges toward the target Boltzmann distribution defined by the energy function in Eq. (4).
@@
We perform a detailed analysis of the proton configurations for $r_{s}$ values larger than 1.41 to gain deeper insight into the molecular phase. Panel (a) of Figure 4 presents a snapshot of the sampled atomic configurations at $r_{s}=1.43$. Hydrogen molecules are identified by pairing the closest hydrogen atoms, with their centers of mass marked by red spheres. Notably, these red spheres do not follow a random liquid arrangement, but instead form an ordered structure reminiscent of a solid. To further investigate these structures, we plot the averaged simulated X-ray diffraction (XRD) spectra for an ensemble of hydrogen configurations at three different $r_{s}$ values in Figure 4 (c). The delta peaks in the XRD spectra indicate that these systems are in a solid state. We quench the atomic configurations by sampling from $p(\bm{X})^{\beta}$ with $\beta=20$ to better characterize the structures. Figure  4 (b) shows the RDF of the original and quenched structures. As the system cools down, the features around $2~{}a_{B}$ and $3~{}a_{B}$ become more prominent. Note that in this work, we use a cubic unit cell setup as in all the liquid studies, which prevents us from examining the detailed crystal structure of the solid state. Nevertheless, our results clearly indicate the emergence of a high-temperature molecular solid at the temperature where PBE-level calculations predict a liquid-liquid phase transition.
@@ IV Conclusion and Discussion
In this work, we observe evidence of a direct transition from atomic liquid to molecular crystal within a temperature regime believed to be relevant to the liquid-liquid transition in the dense hydrogen system. This finding arises from a variational free energy calculation with many-body precision, demonstrating that the flow model is capable of transforming the structureless liquid state of the base proton distribution into a solid-state solution using many-body precision energy. However, since the simulation is performed with a finite box containing 32 hydrogen atoms, finite-size effects may bias both the existence and the structure of the discovered solid states. Nevertheless, this finding aligns with recent discoveries of a possible solid-state phase in similar temperature and pressure regimes obtained through different computational methods and much larger simulation cells [27, 28], although with different detailed characterizations of the solid-state phase. Collectively, these works suggest that the liquid to solid transition may occur at a higher temperature than previously believed, and having precise many body energy beyond DFT is crucial for stabilizing the solid at high temperature. Replacing the expensive wavefunction model with a cheaper yet accurate machine learning force field could allow larger scale calculations, potentially bypassing some of the limitations of this study. Furthermore, nuclear quantum effects can be incorporated into the calculation by generalizing the proton distribution to a variational density matrix, using the neural canonical transformation approach [45, 46].
@@ Acknowledgements.
We thank Guglielmo Mazzola, Yubo Yang, and Qi Zhang for useful discussions. We thank Guglielmo Mazzola for providing the 64 hydrogen atoms reference data shown in Appendix B. This project is supported by the National Natural Science Foundation of China under Grants No. 12122103, No. 92270107, No. T2225018, No. 12188101, No. T2121001 and the Strategic Priority Research Program of Chinese Academy of Sciences under Grants No. XDB0500000 and No. XDB30000000.
Data Availability
The numerical data used to generate all figures in this manuscript are available in the GitHub repository [47].
@@ A.3 Electron wavefunction
Appendix B Benchmarks of the single configuration variational calculations
Table 2: Comparisons of energies of 64H atoms at different temperatures and $r_{s}$ values.
Ha/atom
Ref. 43 This work
1200K12
$r_{s}=1.34$ 12-0.52106(4) 12-0.52311(5)
$r_{s}=1.42$ 12-0.54047(4) 12-0.54257(5)
6000K12
$r_{s}=1.82$ 12-0.55229(4) 12-0.55417(4)
$r_{s}=1.94$ 12-0.56591(4) 12-0.56801(3)
To demonstrate the accuracy of our neural-network wavefunction for single-configuration training, we benchmark it against the results computed using the setup in Ref. 43 for 64 hydrogen atoms under periodic boundary conditions at different densities, as shown in Table 2. At $r_{s}=1.34$ and 1200K, as well as $r_{s}=1.82$ and 6000K, the system is in the atomic phase; at $r_{s}=1.42$ and 1200K, as well as $r_{s}=1.94$ and 6000K, the system is in the molecular phase. The results clearly shows that the neural network wavefunction employed in this paper consistently achieves lower variational energies across all test cases.
Appendix C The DeePMD model
The DeePMD model used in this work is developed via the concurrent learning algorithm implemented in the DP-GEN package [50]. The model’s quality is iteratively enhanced by expanding the training dataset with data that are essential for improving its accuracy in this framework. Each iteration of concurrent learning involves three phases: training, exploration, and DFT calculation.
DFT-calculated energies, forces, and virial tensors serve as labels for training the DeePMD model. The calculations in this work are performed using the ABACUS package [51, 52], employing the Perdew-Burke-Ernzerhof (PBE) generalized gradient approximation [53] for the exchange-correlation functional and a plane-wave basis. The plane-wave energy cutoff is set to 100 Ry, and the k-space is sampled with a grid spacing of 0.15 bohr-1.
Starting with an initial dataset generated from short ab initio molecular dynamics (AIMD) simulations of known hydrogen phases I, II, and III, we trained an ensemble of four DeePMD models using the smooth edition of the DeePMD descriptor [54]. Each model employs a cut-off radius of 6 Å, an embedding network with neuron counts of [25, 50, 100], and a fitting network with neuron counts of [240, 240, 240]. All models share the same training data and hyperparameters but use different random seeds for parameter initialization. Each model is trained for 100,000 steps.
During the exploration phase, isothermal-isobaric (NPT) MD simulations are performed by sampling configurations within a temperature range of 50-3000 K and a pressure range of 10-300 GPa. In the early iterations, simulations primarily focus on relatively low temperatures and pressures, while the later iterations extend to higher temperatures and pressures. Throughout the molecular dynamics simulation trajectories, the maximum standard deviation of the atomic forces—referred to as the model deviation—is used to estimate force-prediction errors. Frames exhibiting a model deviation exceeding 0.5 times the mean absolute magnitude of the predicted forces are flagged as significantly erroneous and potentially indicative of non-physical configurations, such as atomic overlaps. Up to 300 frames are randomly sampled from those exhibiting model deviation values within a candidate range, defined as 0.2 to 0.5 times the mean absolute magnitude of the forces. The energies, forces, and virial tensors associated with these configurations are computed using DFT and incorporated into the training dataset for subsequent iterations. This exploration process gradually enriches the dataset with liquid configurations in addition to the solid ones, enhancing the model’s ability to describe both phases.
As the iterations proceed, the model’s quality gradually improves, reducing the number of configurations whose model deviations are within the candidate range. Once the number of such configurations no longer decreases, we consider the concurrent learning process convergent. In total, 84 concurrent learning iterations were conducted, resulting in 28,537 labeled configurations in the training dataset. A final production DeePMD model was then trained on this dataset for 16,000,000 steps, achieving training accuracies of 0.012 eV/atom for energy, 0.276 eV/Å for forces, and 0.021 eV/atom for the virial tensor. The trained model and the input files can be found in the Github repository https://github.com/fermiflow/SolidHydrogen.
Appendix D Convergence of the proton distribution
Figure 6: Comparison between the proton distribution sampled from the flow model and the target distribution defined by the electronic wavefunction of 32H at $T=1200$ K and $r_{s}=1.44$ (with the minimum value of $-\ln p(\bm{X})$ shifted to zero). (a) Scatter plot comparing the model density and the target density. (b) Histogram of the differences between the two distributions.
To inspect the convergence of the proton distribution parameterized by the flow model (see Eq. 3), we follow the analysis in Ref. 55. Given an energy function $E(\mathbf{X})$, the target proton distribution is defined as $e^{-\beta E(\mathbf{X})}/\mathcal{Z}$, where $\beta=1/k_{B}T$ and $\mathcal{Z}$ is the partition function. At convergence, the flow-based distribution satisfies $p(\mathbf{X})\propto e^{-\beta E(\mathbf{X})}$, which is equivalent to
$\text{ln}p(\mathbf{X})=-\beta E(\mathbf{X})+\text{constant}\,,$ (22)
with the constant related to the unknown partition function. Figure 6 compares distributions computed for the trained flow model and the base distribution, with the constant shift estimated by the average difference between $\beta E(\mathbf{X})$ and $-\text{ln}p(\mathbf{X})$ of the trained flow model. In our calculations, the energy $E(\mathbf{X})$ is evaluated using a trial wavefunction similar to the VMC method, and is therefore subject to statistical noise. We sampled 128 proton configurations for each distribution and generated 3200 electron configurations for each proton configuration to estimate the energy. From the scatter plot shown in panel (a), it is evident that with the trained flow model, most samples lie around the reference line with slope one, whereas the samples from the base distribution exhibit significantly larger deviations and do not show any linear relation. The histograms in panel (b) further demonstrate that the distribution of the trained model shifts toward lower values as a result of the variational optimization of the flow model. Moreover, the distribution becomes more sharply peaked, reflecting a reduction in the variance of the variational free energy.